# Listenable Maps for Audio Classifiers

### v1

@@ Abstract
Despite the impressive performance of deep learning models across diverse tasks, their complexity poses challenges for interpretation. This challenge is particularly evident for audio signals, where conveying interpretations becomes inherently difficult. To address this issue, we introduce Listenable Maps for Audio Classifiers (L-MAC), a posthoc interpretation method that generates faithful and listenable interpretations. L-MAC utilizes a decoder on top of a pretrained classifier to generate binary masks that highlight relevant portions of the input audio. We train the decoder with a special loss that maximizes the confidence of the classifier decision on the masked-in portion of the audio while minimizing the probability of model output for the masked-out portion. Quantitative evaluations on both in-domain and out-of-domain data demonstrate that L-MAC consistently produces more faithful interpretations than several gradient and masking-based methodologies. Furthermore, a user study confirms that, on average, users prefer the interpretations generated by the proposed technique.
@@ 1 Introduction
To summarize, our contributions is the following:
@@ 1.1 Related work
In the audio domain, notable works on interpretability include (Becker et al., 2023), which proposed layer-wise relevance propagation to generate saliency maps over spectrograms. Additionally, (Won et al., ) proposed creating visualizations using attention layer outputs, while (Muckenhirn et al., 2019) suggested using Guided Backpropagation for spectrogram saliency maps. SLIME (Mishra et al., 2017, 2020) proposes to divide the spectrogram into predefined time/frequency regions (akin to the superpixels in LIME (Ribeiro et al., 2016) for images) and determines the feature importance for each region. AudioLIME (Haunschmid et al., 2020; Chowdhury et al., 2021), on the other hand, defines the LIME superpixels as sources extracted from the input audio and determines a saliency score for each source. More recently (Parekh et al., 2022) proposed Listen-to-Interpret (L2I) to learn the classifier relevance for Non-Negative Matrix factorization (NMF) (Lee & Seung, 1999) dictionaries, via a decoder trained to estimate NMF activations. This work is particularly relevant to our method, as we also train a decoder. Consequently, we include detailed comparisons with L2I in our experiments.
@@ 2 Methodology
To generate an interpretation signal that explains these predictions, the latent representations of the classifier are fed into the decoder. The decoder is trained to generate a binary mask , selecting relevant portions of the original spectrogram . Importantly, the decoder is not specifically tuned to the features used by the pretrained classifier as we apply the binary mask to the (linear) spectrogram of the original audio . This feature-agnostic aspect allows L-MAC to maintain listenability. By inheriting the phase from the original audio, we can indeed invert the masked spectrogram and produce a listenable audio signal as an interpretation outcome. To train the decoder, we feed the masked input into the classifier, and compute the mask loss. The objective is to learn a decoder capable of outputting a mask that accurately selects the region of interest in the input spectrogram. The main components of L-MAC are detailed in the following sub-sections.
@@ 2.1 The masking objective
One important aspect of our work is the use of a neural network, specifically the decoder, to estimate the binary mask . This choice is motivated by our observation that neural networks yield more faithful and understandable masks. After parameterizing the interpretation mask with a neural decoder, represented by parameters denoted as , the corresponding optimization objective becomes the following
@@
2.2 Listenable explanations
State-of-the-art audio and speech classifiers often rely on features computed on top of the linear spectrogram, such as FBANKs (mel-spectrograms). These features intelligently compress the frequency axes, often leading to more compact features that enhance performance. However, these features are non-invertible due to the compression applied to the frequency axis.
@@ 3 Experiments
In our experiments, we evaluate the faithfulness and understandability of the generated interpretations. To achieve this, we considered two setups: i) Single-label classification under out-of-domain conditions, and ii) Multi-label classification. We utilized the ESC50 dataset (Piczak, ) which contains 50 environmental sound classes for both setups.
@@ 3.2 Faithfulness Evaluation
| STFTMel | Guided Backprop | 0.00 | 15.66 | 0.00 | 0.05 | 0.06 | 0.39 | 5.47 |
@@
In these experiments, we first train a CNN14 classifier (Kong et al., 2020) on the ESC-50 dataset (Piczak, ) augmented with WHAM! noise, to simulate real-world mixtures. The classifier is trained on folds 1, 2, and 3 and obtains 75% and 78% classification accuracy on folds 5 and 4, respectively. The CNN14 classifier we employed has 12 2D convolutional layers and is pre-trained on the VGG-sound dataset (Chen et al., 2020a) using SimCLR (Chen et al., 2020b). The decoder of L-MAC consists of a series of transposed 2D convolutional layers. Each convolutional layer upsamples the time and frequency axis. The classifier’s representations are fed to the decoder in a U-Net-like fashion to incorporate information at different time-frequency resolutions. Specifically, the decoder takes the four deepest representations of the classifier.
@@
We then freeze the weights of the classifier and train a decoder on the same training set as the classifier (ESC50 + WHAM! noise), as shown in Figure 1. We perform two sets of evaluations on the ESC-50 dataset to validate the robustness of L-MAC in real-world settings. First, we evaluate L-MAC on in-domain data, where the interpreter is tested against data similar to the one in the training set (ESC 50 folds 4 and 5 with WHAM! noise). The results are provided in Table 1. In the second place, we evaluate L-MAC performance on out-of-domain data. We generate out-of-domain data by creating mixtures of samples from folds 4 and 5. The results are provided in Table 2. We report results obtained with STFT and Mel domain spectra for in-domain and out-of-domain data.
@@
To compare L-MAC with the literature, we used several gradient-based methods such as standard saliency maps (Simonyan et al., 2014), SmoothGrad (Smilkov et al., 2017), IntegratedGradients (Sundararajan et al., 2017), GuidedBackProp (Springenberg et al., 2015), and decoder based audio specific method, Listen-to-Interpret (L2I) (Parekh et al., 2022), and we also include SHAP (Lundberg & Lee, 2017). For L2I, we have reported results using relevance thresholds RT=0.2, 0.4, 0.6, and 0.8 (an important hyperparameter for the L2I method). We have used the Captum implementations (Kokhlikyan et al., 2020) for the gradient-based methods and SHAP and adapted the SpeechBrain (Ravanelli et al., 2021) implementation for L2I. For L-MAC, we have obtained results for finetuning strengths . The implementation of our experimental setup is available here111https://anonymous.4open.science/r/audio_interpretability-87B5/recipes/ESC50.
@@
For in-domain data, we observe in Table 1 that L-MAC generally results in better faithfulness scores (AI, AD, AG, FF, Fid-In) compared to the baselines. Since we work with a classifier trained in the Mel domain, we would like to note that we evaluated the methods in cases where we both mask in the STFT domain and in the Mel domain. This is important, as masking in the STFT domain gives the ability to produce listenable interpretations. We observe that for L-MAC, the STFT domain interpretations result in similar faithfulness scores compared to the Mel domain. More specifically, for the in-domain evaluation, we observe that the gradient-based methods such as standard saliency, Smoothgrad, IG, GradCAM, Guided BackProp, and SHAP generally result in less faithful interpretations than the decoder-based methods such as L2I and L-MAC. We see that the most faithful results are obtained without having the finetuning (FT) stage in L-MAC. However, with additional finetuning, we see that L-MAC can produce results that only lose slightly from their faithfulness scores while increasing the understandability of the interpretation, as shown by user preference. We observe that after finetuning, L-MAC can still be more faithful compared to the other baselines we investigate, including our L2I implementation. We also see that in terms of the sparsity metric, L-MAC has better overall numbers than L2I and the other gradient-based methods. In terms of the complexity metric, it is comparable with the gradient-based methods and with L2I with a relevance threshold of 0.2. Note that with larger relevance threshold values, L2I returns less active interpretations.
@@
In Table 2, we show the metrics on out-of-distribution data. In this case, the gradient-based posthoc saliency methods also do not yield very faithful results except for GradCAM. We also observe that L-MAC outperforms L2I even if a severe finetuning strength of is employed. Another observation is that the listenable STFT version of L-MAC yields comparable faithfulness results to the direct Mel domain interpretations yielded by L-MAC. For this data, we also try increasing the cross-correlation threshold (denoted with CCT in Table 2) between the interpretations and the target mask during training (as discussed in Section 2). We see that using a higher cross-correlation threshold generally helps increase the interpretations’ faithfulness for larger values.
@@ 3.3 Qualitative Evaluation
In addition to the showcase recordings for L2I, we have also randomly picked five mixture recordings that we have created and presented these recordings to the users. In total therefore we have presented each user 9 recordings. The average preference for these recordings are shown in Table 3 in the MOS-2 column. Once again, we observe that L-MAC surpasses L2I, and fine-tuning further enhances user preference. The results of this user study for each recording are depicted in the right panel of Figure 2. Some examples of interpretations can be found at our companion website333https://anonymous.4open.science/w/audio_interpretability-87B5/.
@@ 3.4 Sanity Checks
We have also conducted the cascading randomization test from (Adebayo et al., 2020) in order to verify that L-MAC does not generate interpretations that are invariant to the classifier weights, but actually is sensitive to the randomization of the classifier weights. In Figure 3, we showcase this, where we compare the interpretations obtained when the classifier layers are randomized. We see that with more random layers L-MAC interpretations lose their original focus. Whereas for instance with GradCAM interpretations, we observe that the interpretations basically remain insensitive to the classifier weight randomization. We furthermore quantify this finding by calculating the Structural Similarity Index (SSIM) between the interpretations generated for the original model and the interpretations generated for the model with randomized weights (as was done in (Adebayo et al., 2020)). Similar to the visualization of the interpretations in Figure 3, we see that L-MAC interpretations quicly drop in similarity after starting from the 5’th convolutional block (5’th deepest block - note that we start randomizing from the, last layer, and then go down). Whereas, we see that the GradCAM interpretations remain practically unchanged.
@@ 4 Conclusions
This paper introduced a novel approach, called Listenable Maps for Audio Classifiers (L-MAC), to produce the post-hoc interpretations for audio classifiers. L-MAC employs a decoder that leverages the latent representations of the black-box classifier to estimate a binary mask that effectively highlights the portions of audio that triggered the prediction made by the classifier. The application of this mask to the linear spectrogram enables the generating interpretations that are not only listenable but are also applicable for classifiers whose input domains are not linear spectra. We train the decoder using a special composite objective, that minimizes the categorical cross-entropy for masked inputs while maximizing it for masked-out inputs.
@@
Through an extensive experimental evaluation involving quantitative and qualitative assessments, as well as thorough sanity checks, our results demonstrate that L-MAP achieves significantly superior faithfulness and higher quality audio, even when compared to the recently proposed L2I method.

### v2

@@ Abstract
Despite the impressive performance of deep learning models across diverse tasks, their complexity poses challenges for interpretation. This challenge is particularly evident for audio signals, where conveying interpretations becomes inherently difficult. To address this issue, we introduce Listenable Maps for Audio Classifiers (L-MAC), a posthoc interpretation method that generates faithful and listenable interpretations. L-MAC utilizes a decoder on top of a pretrained classifier to generate binary masks that highlight relevant portions of the input audio. We train the decoder with a loss function that maximizes the confidence of the classifier decision on the masked-in portion of the audio while minimizing the probability of model output for the masked-out portion. Quantitative evaluations on both in-domain and out-of-domain data demonstrate that L-MAC consistently produces more faithful interpretations than several gradient and masking-based methodologies. Furthermore, a user study confirms that, on average, users prefer the interpretations generated by the proposed technique.
@@ 1 Introduction
To summarize, our contributions are the following:
@@ 1.1 Related Works
In the audio domain, notable works on interpretability include (Becker et al., 2023), which proposed layer-wise relevance propagation to generate saliency maps over spectrograms. Another noteworthy works include (Trinh et al., 2018; Kavaki & Mandel, 2020) where authors learn to identify important parts of the input spectrogram by masking additive white noise, within the context of automatic speech recognition. Additionally, (Won et al., 2019) proposed creating visualizations using attention layer outputs, while (Muckenhirn et al., 2019) suggested using Guided Backpropagation for spectrogram saliency maps. SLIME (Mishra et al., 2017, 2020) proposes to divide the spectrogram into predefined time/frequency regions (akin to the superpixels in LIME (Ribeiro et al., 2016) for images) and determines the feature importance for each region. AudioLIME (Haunschmid et al., 2020; Chowdhury et al., 2021), on the other hand, defines the LIME superpixels as sources extracted from the input audio and determines a saliency score for each source. More recently (Parekh et al., 2022) proposed Listen-to-Interpret (L2I) to learn the classifier relevance for Non-Negative Matrix factorization (NMF) (Lee & Seung, 1999) dictionaries, via a decoder trained to estimate NMF activations. This work is particularly relevant to our method, as we also train a decoder. Consequently, we include detailed comparisons with L2I in our experiments.
@@ 2 Methodology
To generate an interpretation signal that explains these predictions, the latent representations of the classifier are fed into the decoder, as illustrated in Figure 1. A more detailed diagram of how the classifier representations are used is shown in Figure 5 in Appendix A. The decoder is trained to generate a binary mask , selecting relevant portions of the original spectrogram . Importantly, the decoder is not specifically tuned to the features used by the pretrained classifier as we apply the binary mask to the (linear) spectrogram of the original audio . This feature-agnostic aspect allows L-MAC to maintain listenability. By inheriting the phase from the original audio, we can indeed invert the masked spectrogram and produce a listenable audio signal as an interpretation outcome. To train the decoder, we feed the masked input into the classifier, and compute the mask loss. The objective is to learn a decoder capable of outputting a mask that accurately selects the region of interest in the input spectrogram. The main components of L-MAC are detailed in the following sub-sections.
@@ 2.1 The Masking Objective
One important aspect of our work is the use of a neural network, specifically the decoder, to estimate the binary mask . This choice is motivated by our observation that neural networks yield more faithful and understandable masks. After parameterizing the interpretation mask with a neural decoder, represented by parameters denoted as , the corresponding optimization objective becomes the following,
@@
2.2 Producing Listenable Explanations
State-of-the-art audio and speech classifiers often rely on features computed on top of the linear spectrogram, such as FBANKs (mel-spectrograms). These features intelligently compress the frequency axis, often leading to more compact features that enhance performance. However, these features are non-invertible due to the compression applied to the frequency axis.
@@ 3 Experiments
In our experiments, we evaluate the faithfulness and understandability of the generated interpretations. To achieve this, we considered two setups: i) Classification under in-domain conditions, and ii) Classification under out-of-domain conditions. We utilized the ESC50 dataset (Piczak, ) which contains 50 environmental sound classes for both setups. We also provide additional results on the UrbanSound8k dataset (Salamon et al., 2014) in Appendix B.
@@ 3.2 Faithfulness Evaluation
| STFT-Mel | Guided Backprop | 0.00 | 15.66 | 0.00 | 0.05 | 0.06 | 0.39 | 5.47 |
@@
In these experiments, we first train a CNN14 classifier (Kong et al., 2020) on the ESC-50 dataset (Piczak, ) augmented with WHAM! noise, to simulate real-world mixtures. The classifier is trained on folds 1, 2, and 3 and obtains 75% and 78% classification accuracy on folds 5 and 4, respectively. The CNN14 classifier we employed has 12 2D convolutional layers and is pre-trained on the VGG-sound dataset (Chen et al., 2020a) using SimCLR (Chen et al., 2020b). The decoder of L-MAC consists of a series of transposed 2D convolutional layers. Each convolutional layer upsamples the time and frequency axis. The classifier’s representations are fed to the decoder in a U-Net-like fashion to incorporate information at different time-frequency resolutions (as shown in Figure 5 in Appendix A). Specifically, the decoder takes the four deepest representations of the classifier.
@@
We then freeze the weights of the classifier and train a decoder on the same training set as the classifier (ESC50 + WHAM! noise), as shown in Figure 1. We perform two sets of evaluations on the ESC-50 dataset to validate the robustness of L-MAC in real-world settings. First, we evaluate L-MAC on in-domain data, where the interpreter is tested against data similar to the one in the training set (ESC 50 folds 4 and 5 with WHAM! noise). The results are provided in Table 1. In the second place, we evaluate L-MAC performance on out-of-domain data. We generate out-of-domain data by creating mixtures of samples from folds 4 and 5. The results are provided in Table 2. We report results obtained with STFT and Mel domain spectra for in-domain and out-of-domain data. We have also reported additional results on the UrbanSound8k dataset (Salamon et al., 2014) in Table 2, in Appendix B.
@@
To compare L-MAC with the literature, we used several gradient-based methods such as standard saliency maps (Simonyan et al., 2014), SmoothGrad (Smilkov et al., 2017), IntegratedGradients (Sundararajan et al., 2017), GuidedBackProp (Springenberg et al., 2015), and decoder based audio specific method, Listen-to-Interpret (L2I) (Parekh et al., 2022), and we also include SHAP (Lundberg & Lee, 2017). For L2I, we have reported results using relevance thresholds RT=0.2, 0.4, 0.6, and 0.8 (an important hyperparameter for the L2I method). We have used the Captum implementations (Kokhlikyan et al., 2020) for the gradient-based methods and SHAP and adapted the SpeechBrain (Ravanelli et al., 2021) implementation for L2I. For L-MAC, we have obtained results for finetuning strengths . The implementation of our experimental setup is accessible through the companion website.111https://fpaissan.github.io/lmac.
@@
For in-domain data, we observe in Table 1 that L-MAC generally results in better faithfulness scores (AI, AD, AG, FF, Fid-In) compared to the baselines. Since we work with a classifier trained in the Mel domain, we would like to note that we evaluated the methods in cases where we both mask in the STFT domain (denoted with STFT-Mel in Tables 1, 2) and in the Mel domain (denoted with Mel in Tables 1, 2). This is important, as masking in the STFT domain gives the ability to produce listenable interpretations. We observe that for L-MAC, the STFT domain interpretations result in similar faithfulness scores compared to the Mel domain. More specifically, for the in-domain evaluation, we observe that the gradient-based methods such as standard saliency, Smoothgrad, IG, GradCAM, Guided BackProp, and SHAP generally result in less faithful interpretations than the decoder-based methods such as L2I and L-MAC. We see that the most faithful results are obtained without having the finetuning (FT) stage in L-MAC. However, with additional finetuning, we see that L-MAC can produce results that only lose slightly from their faithfulness scores while increasing the understandability of the interpretation, as shown by user preference. We observe that after finetuning, L-MAC can still be more faithful compared to the other baselines we investigate, including our L2I implementation. We also see that in terms of the sparsity metric, L-MAC has better overall numbers than L2I and the other gradient-based methods. In terms of the complexity metric, it is comparable with the gradient-based methods and with L2I with a relevance threshold of 0.2. Note that with larger relevance threshold values, L2I returns less active interpretations.
@@
In Table 2, we show the metrics on out-of-distribution data. In this case, the gradient-based posthoc saliency methods also do not yield very faithful results except for GradCAM. We also observe that L-MAC outperforms L2I even if a severe finetuning strength of is employed. Another observation is that the listenable STFT version of L-MAC yields comparable faithfulness results to the direct Mel domain interpretations yielded by L-MAC. For this data, we also try increasing the cross-correlation threshold (denoted with CCT in Table 2) between the interpretations and the target mask during training (as discussed in Section 2). We see that using a larger CCT generally helps increase the interpretations’ faithfulness for larger values.
@@ 3.3 Qualitative Evaluation
In addition to the showcase recordings for L2I, we have also randomly picked five mixture recordings that we have created and presented these recordings to the users. In total therefore we have presented each user 9 recordings. The average preference for these recordings are shown in Table 3 in the MOS-2 column. Once again, we observe that L-MAC surpasses L2I, and fine-tuning further enhances user preference. The results of this user study for each recording are depicted in the right panel of Figure 2. Some examples of interpretations can be found at our companion website333https://fpaissan.github.io/lmac/.
@@ 3.4 Sanity Checks
We have also conducted the cascading randomization test from (Adebayo et al., 2020) in order to verify that L-MAC does not generate interpretations that are invariant to the classifier weights, but actually is sensitive to the randomization of the classifier weights. In Figure 3, we showcase this, where we compare the interpretations obtained when the classifier layers are randomized. We see that, as expected, with more random layers L-MAC interpretations lose their original focus. Whereas for instance with GradCAM interpretations, we observe that the interpretations basically remain insensitive to the classifier weight randomization. We furthermore quantify this finding by calculating the Structural Similarity Index (SSIM) between the interpretations generated for the original model and the interpretations generated for the model with randomized weights (as was done in (Adebayo et al., 2020)). Similar to the visualization of the interpretations in Figure 3, we see that L-MAC interpretations quicly drop in similarity after starting from the 5’th convolutional block (5’th deepest block - note that we start randomizing from the, last layer, and then go down). Whereas, we see that the GradCAM interpretations remain practically unchanged.
| Metric | AI () | AD () | AG () | FF () | Fid-In () | SPS () | COMP () | MM |
| Classification on ESC50, White Noise Contamination, 38.6% accuracy | ||||||||
| Saliency | 0.25 | 26.31 | 0.02 | 0.05 | 0.06 | 0.79 | 10.92 | 0.016 |
| Smoothgrad | 0.00 | 26.37 | 0.00 | 0.04 | 0.09 | 0.84 | 10.62 | 0.01 |
| IG | 0.75 | 25.60 | 0.56 | 0.10 | 0.21 | 0.82 | 10.65 | 0.01 |
| L2I @ 0.2 | 0.00 | 19.41 | 0.21 | 0.11 | 0.04 | 36.62 | 7.32 | 0.12 |
| GradCAM | 8.87 | 20.88 | 1.24 | 0.28 | 0.51 | 0.69 | 11.25 | 0.18 |
| Guided GradCAM | 0.50 | 26.23 | 0.05 | 0.07 | 0.11 | 0.91 | 10.14 | 0.01 |
| Guided Backprop | 0.25 | 26.30 | 0.02 | 0.05 | 0.07 | 0.79 | 10.92 | 0.02 |
| SHAP | 0.12 | 26.34 | 0.001 | 0.05 | 0.12 | 0.86 | 10.40 | 0.004 |
| L-MAC (ours) | 83.62 | 1.50 | 56.12 | 0.33 | 0.86 | 0.92 | 10.03 | 0.06 |
| All-ones baseline | 0 | 0 | 0 | 0.34 | 1 | N.A. | N.A. | 1 |
| Classification on ESC50, LJSpeech Contamination, 79.3% accuracy | ||||||||
| Saliency | 0.87 | 26.00 | 0.20 | 0.06 | 0.11 | 0.75 | 11.10 | 0.02 |
| Smoothgrad | 0.50 | 26.14 | 0.11 | 0.05 | 0.13 | 0.79 | 10.91 | 0.08 |
| IG | 0.37 | 25.70 | 0.01 | 0.11 | 0.25 | 0.87 | 10.14 | 0.00 |
| L2I @ 0.2 | 1.75 | 29.49 | 0.27 | 0.15 | 0.18 | 0.79 | 9.56 | 0.16 |
| GradCAM | 20.37 | 13.49 | 2.63 | 0.28 | 0.73 | 0.66 | 11.33 | 0.22 |
| Guided GradCAM | 0.25 | 26.10 | 0.09 | 0.06 | 0.11 | 0.88 | 10.30 | 0.01 |
| Guided Backprop | 0.87 | 26.01 | 0.20 | 0.05 | 0.11 | 0.75 | 11.10 | 0.02 |
| SHAP | 0.00 | 26.14 | 0.00 | 0.06 | 0.16 | 0.79 | 10.81 | 0.01 |
| L-MAC (ours) | 70.75 | 2.73 | 39.64 | 0.33 | 0.83 | 0.93 | 9.70 | 0.05 |
| All-ones baseline | 0 | 0 | 0 | 0.35 | 1 | N/A | N/A | 1 |
3.5 Additional Results on Out-of-Domain Data
In addition to the Out-of-Domain experiments conducted in Section 2, we have also tested L-MAC on audio samples corrupted with white noise and speech. We have created 3dB Signal-to-noise ratio mixtures, and used samples from the LJSpeech (Ito & Johnson, 2017) dataset for speech. In Table 4, we show these additional results on ESC50 dataset. In Appendix B with Table 5, we provide the results for this experimental setup applied on the UrbanSound8K benchmark. We observe that L-MAC is able to obtain better results in terms of quantitative faithfulness metrics such as AI, AD, AG, FF, and Fid-In. In this table we also report the mask-mean to indicate the size of the obtained mask (denoted with MM). We observe that even though the mask area of L-MAC is significantly smaller than GradCAM for instance, it is able to obtain considerably better metrics that measure faithfulness. Moreover, we have also calculated the scores obtained with an all-zeros mask. We observe that the FF score obtained with L-MAC is very similar to the all-ones mask which removes the entire spectrogram during the score computation, and we also observe that the mask-mean of L-MAC is significantly smaller than 1. This indicates that the masked-out portion of the L-MAC interpretations are not very significant for the classifier.
@@ 4 Conclusions
This paper introduced a novel approach, called Listenable Maps for Audio Classifiers (L-MAC), that produces post-hoc interpretations for audio classifiers. L-MAC employs a decoder that uses the latent representations of the black-box classifier to estimate a binary mask that effectively highlights the portions of the input audio that triggered the prediction made by the classifier. The application of this mask to the linear-frequency-scale spectrogram enables generating interpretations that are listenable. We would like to note that our methodology is also applicable for classifiers whose input domains are not linear-frequency-scale spectra. We train the decoder using an objective that promotes faithfulness to the classifier decisions, that minimizes the categorical cross-entropy for masked inputs while maximizing it for masked-out inputs.
@@
Through an extensive experimental evaluation involving quantitative and qualitative assessments, as well as through sanity checks, our results demonstrate that L-MAC achieves significantly superior faithfulness metrics and user preference, compared to various baselines.
Impact statement
This paper proposes a methodology to improve neural network interpretability. We do not foresee a direct, socially negative impact. On the contrary, the proposed methodology may facilitate socially beneficial applications of machine learning-based audio processing by enhancing the human trust in neural network decisions.
Acknowledgements
We thank Geraldin Nanfack and Eugene Belilovsky for insightful discussions. This research was enabled in part by support provided by Calcul Québec and the Digital Research Alliance of Canada.
Appendix A Decoder structure illustration
In Figure 5 we show the architecture of the decoder employed in our experiments.
Appendix B Results on UrbanSound8k Dataset
We provide additional results for the In-Domain and Out-of-Domain experiments on the UrbanSound8k dataset (Salamon et al., 2014), in addition to the experiments conducted on the ESC50 dataset in the mainbody of the paper. In Table 5, we first report the results obtained on the In-Domain experiments, and then for the Out-of-Domain experiments where we contaminate the samples with other samples from the US8k dataset, White Noise, and speech (the same experimental setup as the setup on the ESC50). We observe that on both the ID experiment and on different OOD experiments, L-MAC is able to outperform or obtain very similar results compared to the baselines. We also observe that L-MAC masks are in general smaller compared to the ones obtained by comparably performing models such as GradCAM or L2I, which suggests that L-MAC is able to obtain masks that are not trivial.
| Metric | AI () | AD () | AG () | FF () | Fid-In () | SPS () | COMP () | MM |
| Classification on US8k, WHAM! Noise Contamination (ID experiment), 82.8% accuracy | ||||||||
| Saliency | 0.40 | 23.86 | 0.26 | 0.072 | 0.12 | 0.30 | 5.69 | 0.02 |
| Smoothgrad | 0.00 | 26.86 | 0.00 | 0.04 | 0.04 | 0.30 | 5.69 | 0.01 |
| IG | 0.25 | 23.70 | 0.15 | 0.10 | 0.12 | 0.40 | 5.24 | 0.005 |
| L2I @ 0.2 | 0.55 | 19.59 | 0.28 | 0.17 | 0.18 | 0.24 | 5.67 | 0.21 |
| GradCAM | 7.68 | 6.94 | 4.04 | 0.31 | 0.39 | 0.13 | 5.96 | 0.29 |
| Guided GradCAM | 0.40 | 24.29 | 0.26 | 0.06 | 0.10 | 0.33 | 5.58 | 0.01 |
| Guided Backprop | 0.40 | 23.86 | 0.26 | 0.07 | 0.12 | 0.30 | 5.69 | 0.02 |
| SHAP | 0.10 | 26.32 | 0.03 | 0.05 | 0.06 | 0.33 | 5.59 | 0.01 |
| L-MAC (ours) | 19.09 | 2.51 | 13.57 | 0.40 | 0.46 | 0.32 | 5.55 | 0.14 |
| All-ones baseline | 0 | 0 | 0 | 0.42 | 1 | N/A | N/A | 1 |
| Classification on US8k, US8k Contamination, 82.6% accuracy | ||||||||
| Saliency | 0.40 | 48.24 | 0.23 | 0.10 | 0.24 | 0.61 | 11.35 | 0.056 |
| Smoothgrad | 0.00 | 54.65 | 0.00 | 0.06 | 0.09 | 0.60 | 11.40 | 0.023 |
| IG | 0.51 | 46.31 | 0.20 | 0.15 | 0.30 | 0.80 | 10.39 | 0.01 |
| L2I @ 20 | 2.88 | 27.84 | 2.01 | 0.27 | 0.38 | 0.50 | 11.05 | 0.37 |
| GradCAM | 14.65 | 9.71 | 6.42 | 0.61 | 0.86 | 0.18 | 12.02 | 0.60 |
| Guided GradCAM | 0.40 | 49.61 | 0.19 | 0.10 | 0.21 | 0.66 | 11.17 | 0.03 |
| Guided Backprop | 0.40 | 48.24 | 0.23 | 0.10 | 0.24 | 0.61 | 11.35 | 0.04 |
| SHAP | 0.05 | 53.66 | 0.01 | 0.07 | 0.12 | 0.67 | 11.17 | 0.01 |
| L-MAC (ours) | 27.12 | 9.58 | 20.04 | 0.77 | 0.85 | 0.71 | 10.75 | 0.24 |
| All-ones baseline | 0 | 0 | 0 | 0.86 | 1 | N/A | N/A | 1 |
| Classification on US8k, White Noise Contamination, 48.5% accuracy | ||||||||
| Saliency | 0.56 | 48.42 | 0.16 | 0.13 | 0.14 | 0.57 | 11.50 | 0.04 |
| Smoothgrad | 0.00 | 48.96 | 0.00 | 0.07 | 0.10 | 0.66 | 11.24 | 0.02 |
| IG | 0.85 | 47.92 | 0.30 | 0.15 | 0.16 | 0.63 | 11.21 | 0.02 |
| L2I @ 0.2 | 27.45 | 32.73 | 23.45 | 0.25 | 0.42 | 0.32 | 11.81 | 0.41 |
| GradCAM | 37.37 | 16.69 | 31.29 | 0.45 | 0.69 | 0.24 | 11.95 | 0.61 |
| Guided GradCAM | 0.51 | 48.26 | 0.27 | 0.12 | 0.14 | 0.62 | 11.31 | 0.03 |
| Guided Backprop | 0.56 | 48.41 | 0.16 | 0.13 | 0.13 | 0.59 | 11.44 | 0.04 |
| SHAP | 0.15 | 48.51 | 0.02 | 0.09 | 0.12 | 0.72 | 10.98 | 0.01 |
| L-MAC (ours) | 31.62 | 16.31 | 22.4 | 0.81 | 0.71 | 0.39 | 11.72 | 0.39 |
| All-ones baseline | 0 | 0 | 0 | 0.74 | 1 | N/A | N/A | 0 |
| Classification on US8k, LJSpeech Contamination, 88.7% accuracy | ||||||||
| Saliency | 0.30 | 50.60 | 0.11 | 0.08 | 0.25 | 0.58 | 11.44 | 0.04 |
| Smoothgrad | 0.00 | 52.68 | 0.00 | 0.05 | 0.19 | 0.61 | 11.39 | 0.02 |
| IG | 0.35 | 50.13 | 0.18 | 0.10 | 0.26 | 0.79 | 10.43 | 0.01 |
| L2I @ 0.20 | 1.21 | 41.91 | 0.72 | 0.23 | 0.27 | 0.55 | 10.50 | 0.23 |
| GradCAM | 13.29 | 7.16 | 6.20 | 0.53 | 0.91 | 0.25 | 11.93 | 0.61 |
| Guided GradCAM | 0.30 | 51.18 | 0.14 | 0.08 | 0.23 | 0.64 | 11.27 | 0.03 |
| Guided Backprop | 0.30 | 50.6 | 0.11 | 0.09 | 0.25 | 0.58 | 11.44 | 0.04 |
| SHAP | 0.05 | 52.15 | 0.01 | 0.06 | 0.18 | 0.67 | 11.16 | 0.01 |
| L-MAC (ours) | 18.18 | 9.91 | 11.28 | 0.90 | 0.86 | 0.69 | 10.91 | 0.26 |
| All-ones baseline | 0 | 0 | 0 | 0.83 | 1 | N/A | N/A | 0 |